OpenStudy (ayonnaleflore):

will fan and medal Solve 2x^2 = 72. {±18} {±36} {±6} {±8}

rishavraj (rishavraj):

see u need to isolate the variable 'x' so first step would be dividing on both sides by 2

OpenStudy (ayonnaleflore):

so x^2 = 36

rishavraj (rishavraj):

now u know wht to do??

rishavraj (rishavraj):

see if a^2 = x a = sqrt{x}

OpenStudy (ayonnaleflore):

take the square root of 36 which is 6

rishavraj (rishavraj):

u see 6^2 = 36 and also {-6}^2 = 36

OpenStudy (ayonnaleflore):

\[\pm6\] would be the answer

rishavraj (rishavraj):

yeah good:))

OpenStudy (ayonnaleflore):

ok thank you ..... so all problems like that are solved by the same way for the mot part

rishavraj (rishavraj):

yup u need to isolate the variable :)

OpenStudy (ayonnaleflore):

ok thank you so much

rishavraj (rishavraj):

u welcome :))